Suppose Ghana produces cocoa and timber, and the following hypothetical data is provided for 2016 and 2017.

2016 2017
Good Quantity Price Quantity Price
Cocoa 1000 2 1100 3
Timber 500 10 525 8

Using 2016 as a base year, real GDP in 2017 is ...
[1] 3125
[2] 7500
[3] 5250
[4] 7450
1
Expert's answer
2019-09-23T09:05:39-0400

[4] 7450


Real GDP is equal to the sum of the base year price×current year quantity of all the goods:

2016: (1000×2) + (500×10) = + 2,000 + 5,000 = $ 7,000.

2017: (1100×2) + (525×10) = + 2,200 + 5,250 = $ 7,450.